Coastal Computer operates two retail outlets in Oakview, one on Main Street and the other in Lakeland Mall. The stores share the use of a central accounting department. The cost of the accounting department for last year was $220,000. Following are the operating results for the two stores for the year.
  

Main Street Lakeland Mall
Sales revenues $ 1,981,560 $ 2,736,440
Number of computers sold 2,262 1,638

  

Required:

a. Allocate the cost of the central accounting department to the two stores based on:

  

1. Number of computers sold. (Do not round intermediate calculations.)

2. Store revenue. (Do not round intermediate calculations.)
Answer #1

Solution:

1.) Number of Computers Sold:
Main Street =$220,000/(2,262+1,638)*2,262 = $127,600
Lakwland Mall = $220,000/(2,262+1,638)*1,638 = $92,400

2.) Store Revenue:
Main Street = 220,000/(1,981,560+2,736,440)*1,981,560 = $92,400
Lakwland Mall = 220,000/(1,981,560+2,736,440)*2,736,440 = $127,600